OpenStudy (anonymous):

7m+n=10 m-3n=14 what is the solution ordered pair?

OpenStudy (anonymous):

7m+n = 10....................(1) m-3n=14 ......................(2) from (1) 21m+3n=30....................(3) add (2) and (3) 22m = 44 m = 2 sub m in (2): 2-3n=14 3n = -12 n = -4 solution is 2,-4

OpenStudy (anonymous):

I am having a hard time with these for some reason. thank you